OpenStudy (anonymous):

can you solve the equation -7x+13=-5x-19

OpenStudy (richyw):

you want to isolate x. so add 5x to each side.

OpenStudy (nottim):

Isolate for x. Keep all values with x on right side, all values without x on left side.

OpenStudy (richyw):

\[-2x+13=-19\]understand that?

OpenStudy (anonymous):

-7x+13=-5x-19 -7x+5x=-19-13 -2x=-32 if you divide both sides by -2 you get x=16

OpenStudy (anonymous):

thanks!!!